OpenStudy (anonymous):

What are the next four terms in the following arithmetic sequence? -7, -2, 3, 8, .

OpenStudy (anonymous):

just keep adding on five and you'll get the next four terms

OpenStudy (anonymous):

example? and can you show your work

OpenStudy (anonymous):

13,18,23,28?

OpenStudy (anonymous):

-7+5=-2 -2+5=3 3+5=8 8+5=13 13+5=18 18+5=23 23+5=28

OpenStudy (anonymous):

\[(8-3=5)~~~~~(3-(-2)=5)~~~~~~(-2-(-7)=5)\]

OpenStudy (anonymous):

Take one of terms and subtract it from the previous one if its *arithmetic* geometric is diff, but anyways thsi number (5 0 is what we add to get the next 4 terms like @megakiller7 explained, to get the previous four terms you subtract 5

OpenStudy (anonymous):

so does that mean the last four numbers are 13, 18, 23, and 28? @doulikepiecauseidont

OpenStudy (anonymous):

You got it! great job!